in order to find out the density of a small rock. eddie dropped the rock into a graduated cylinder fill with 52.5 ml of water. the level of the water went up to 68.1ml. eddie then paced therock on a triple beambalance, and its mass was 27.69g what is the volume of the rock? what is the density of the rock?
1
Expert's answer
2020-03-05T07:56:56-0500

The volume of the rock equels the difference between the initial and final volumes of water where the rock was placed:

V(rock) = V(final) - V (initial) = 68.1 ml - 52.5 ml = 15.6 ml

The density of the rock can be found from the following equation:

d(rock) = m(rock) / V(rock) = 27.69 g / 15.6 ml = 1.78 g/ml


Answer: 15.6l and 1.78 g/ml